OpenStudy (anonymous):

Find an equation for the nth term of a geometric sequence where the second and fifth terms are -21 and 567, respectively.

OpenStudy (anonymous):

I know An=A1+(n−1)d, but I don't know how to find difference or the first term

OpenStudy (anonymous):

no, for geometric series \[r=\frac{ a _{n} }{ a _{n-1} }\] \[t _{n}=a r ^{n-1}\]

OpenStudy (anonymous):

Oh, okay then i really have no clue what I'm doing. I'm trying to read the online lesson but it makes no sense to me, could you explain?

OpenStudy (anonymous):

put n=2, n=5 divide and find r find a find \[t _{n}\]

OpenStudy (anonymous):

to start with let a be the first term and r be common ratio of the given G.P

OpenStudy (anonymous):

so tn=(-21)r (2)−1? I dont really understand

OpenStudy (anonymous):

\[t _{2}=ar ^{2-1}=ar=-21\] \[t _{5}=?\]

OpenStudy (anonymous):

oh, okay.. t5= -21

OpenStudy (anonymous):

no,\[t _{2}=ar=-21 ...(1)\] \[t _{5}=ar ^{5-1}=567\] \[ar^4=567 ...(2)\] divide (2) by (1)

OpenStudy (anonymous):

oh sorry, i meant t3=-21 so ar4= 1,134?? i feel im not understanding

OpenStudy (anonymous):

\[\frac{ ar^4 }{ ar }=\frac{ 567 }{ -21 }\] \[r^3=-27=\left( -3 \right)^3,r=-3\] ar=-21 a*(-3)=-21 \[a=\frac{ -21 }{ -3 }=7\] \[t _{n}=?\]

OpenStudy (anonymous):

plug the values of a and r and find \[t _{n}\]

OpenStudy (anonymous):

tn= (7)(-3)^n−1

OpenStudy (anonymous):

idk why its doing that, i put n-1

OpenStudy (anonymous):

\[t _{n}=7\left( -3 \right)^{n-1}\]

OpenStudy (anonymous):

i got it! thank you so much for taking your time!(:

OpenStudy (anonymous):

i am leaving and good night.